if a sample of caproic acid was found to have 0.140 g carbon, 0.0234 g hydrogen and 0.0622 g of oxygen, what is its empirical formula? if the molar mass of the caproic acid is 116 g/mol, what is it molecular formula


1
Expert's answer
2021-04-22T03:02:58-0400

The number of elements in the formula equals:

n(C) = 0.140 g / 12 g/mol × 1 mol = 0.0117

n(H) = 0.0234 g / 1 g/mol × 1 mol = 0.0234

n(O) = 0.0622 g / 16 g/mol × 1 mol = 0.0039

From here:

n(C) = 3

n(H) = 6

n(O) = 1

As a result, empirical formula is C3H6O.

If the molar mass of the caproic acid is 116 g/mol, the molecular formula is:

Mr(C3H6O)x = Mr(caproic acid)

x = Mr(caproic acid) / Mr(C3H6O) = 116 g/mol / (3 × 12g/mol + 6 × 1 g/mol + 1 × 16 g/mol) = 116 g/mol / 58 g/mol = 2

As a result, the molecular formula is:

(C3H6O)2 or C6H12O2


Answer: C3H6O and C6H12O2